OpenStudy (theghostlycreeper):

What is the value of c? 14=2c−6+3c14=2c−6+3c

alones (alones):

This equation seems confusing..is there any exponents or something..?

alones (alones):

Or maybe can you take a screenshot.

OpenStudy (theghostlycreeper):

thats the only thing

alones (alones):

Unless you mean \[\ 14=2c−6+3c\] \[\ 14=2c−6+3c \]

OpenStudy (theghostlycreeper):

yes

alones (alones):

Oh you repeated it..nvm than

alones (alones):

waitt is it two or one?

OpenStudy (theghostlycreeper):

one

alones (alones):

Alright let's start from simplifying..you should end up with \(\ 14=5c−6\) Than flip the equation \(5c−6=14\) and add \(6\) to both sides \[\ 5c−6+6=14+6\] \[\ 5c=20\]

alones (alones):

Than divide by \(\ 5\) \[\frac{ 5c }{ 5 }=\frac{ 20 }{ 5 }\] Your final answer would be \[\ c=4\]

KjSaif (kjsaif):

@AloneS is correct